# Study of $1^{--}$ P wave charmoniumlike and bottomoniumlike tetraquark spectroscopy

### v2

@@ IV Summary
The PDG established state around 4.23 GeV may split into two distinct states. In the assignment, the Y(4230), observed in the $\pi^{+}\pi^{-}J/\psi$ and $K^{+}K^{-}J/\psi$ channels, is assigned as E1 state of an S=0 $1^{--}$ P-wave tetraquark. The Y’(4230), observed in the $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$ channels, is assigned as E2 state of an S=0 $1^{--}$ P-wave tetraquark.
The PDG established state around 4.36 GeV may also split into three distinct states. In the assignment, the Y(4340) observed only in the $\pi^{+}\pi^{-}J/\psi$ process, is assigned as E3 state of an S=0 $1^{--}$ tetraquark, which can be paired with the E1 Y(4230) state. The Y(4380), observed in the $\pi^{+}\pi^{-}J/\psi$ channel, is assigned as E4 state of an S=0 $1^{--}$ tetraquark, while the Y’(4380), observed in the $\pi^{+}\pi^{-}h_{c}$ channel, is assigned as E5 state of an S=0 $1^{--}$ tetraquark.
The PDG established state Y(4660), observed in the $\pi^{+}\pi^{-}J/\psi$ channel and grouped with Y(4230), Y(4340) and Y(4380) from the same process, is assigned as the E15 S=0 $1^{--}$ tetraquark state.
The Y(4484), observed in the $K^{+}K^{-}J/\psi$ channel by BESIII, is grouped with Y(4230) from the $\pi^{+}\pi^{-}J/\psi$ channel and tentatively assigned as the E8 S=0 $1^{--}$ tetraquark state.
The Y(4544), newly observed in the $\omega\chi_{c1}$ channel by BESIII, is grouped with Y’(4230) from the $\omega\chi_{c0}$ channel and tentatively assigned as E11 S=0 $1^{--}$ tetraquark state.
@@ Appendix A Tetraquark spatial wave function
×$\times\psi_{n_{\chi_{1}}l_{\chi_{1}}}(\vec{\chi}_{1}\,)\otimes\psi_{n_{\chi_{2}}l_{\chi_{2}}}(\vec{\chi}_{2}\,)\otimes\psi_{n_{\chi_{3}}l_{\chi_{3}}}(\vec{\chi}_{3})$
where $\psi_{n_{\chi_{i}}l_{\chi_{i}}}$ are harmonic oscillator wave functions. The sum ${n_{\chi_{i}},l_{\chi_{i}}}$ is over $n_{\chi_{1}},n_{\chi_{2}},n_{\chi_{3}},l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}$. $N$ and $L$ are respectively the total principle quantum number and total angular momentum, and $L=l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$ for P wave tetraquark and $N=(2n_{\chi_{1}}+l_{\chi_{1}})+(2n_{\chi_{2}}+l_{\chi_{2}})+(2n_{\chi_{3}}+l_{\chi_{3}})$.
@@
Appendix B Cross term between different configuration from spin-spin interaction
Table 9: Expectation values of the color-spin operator for the S=0 tetraquark state. $<\psi^{k1}_{c}\psi^{h1}_{s}|\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}|\psi^{k2}_{c}\psi^{h2}_{s}>$
@@ Appendix B Cross term between different configuration from spin-spin interaction
Table 10: Expectation values of the color-spin operator for S=2 tetraquark state. $<\psi^{k1}_{c}\psi^{h1}_{s}|\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}|\psi^{k2}_{c}\psi^{h2}_{s}>$
@@
The color-spin operator $\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}$ in the Hamiltonian in Eq. (II.2), along with the mass-dependent factor in the spin-spin interaction,
@@
mixes different color-spin configurations, leading to nonzero cross terms for charmonium-like and bottomonium-like tetraquarks.

### v3

@@ IV Summary
The PDG established state around 4.23 GeV may be treated as a single state at this stage. In the assignment, the Y(4230), observed in the $\pi^{+}\pi^{-}J/\psi$, $K^{+}K^{-}J/\psi$, $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$ channels, is assigned as S=0 $1^{--}$ tetraquark state.
The PDG established state around 4.36 GeV may split into four distinct states. In the assignment, the Y(4300) and Y(4340) observed only in the $\pi^{+}\pi^{-}J/\psi$ process, are assigned as $1^{--}$ tetraquark state respectively. The Y(4390), observed in the $\pi^{+}\pi^{-}h_{c}$ channel, is assigned as a $1^{--}$ tetraquark state with a large decay ratio to $\omega\chi_{cJ}$, while the Y’(4390), observed in the $\pi^{+}\pi^{-}J/\psi$ channel, is assigned as a tetraquark state with a small decay ratio to $\omega\chi_{cJ}$. The Y(4390) could potentially be observed in the $\omega\chi_{c0}$ channel, which may motivate future experimental searches.
The PDG established state Y(4660), observed in the $\pi^{+}\pi^{-}J/\psi$ channel and grouped with Y’(4390) from the same process, is assigned as the a $1^{--}$ tetraquark state with a small decay ratio to $\omega\chi_{cJ}$.
The Y(4484), observed in the $K^{+}K^{-}J/\psi$ channel by BESIII, and the Y(4544), newly observed in the $\omega\chi_{c1}$ channel by BESIII, are tentatively assigned as $1^{--}$ tetraquark states due to mass agreement, respectively.
@@ Appendix A Tetraquark spatial wave function
$\times\psi_{n_{\chi_{1}}l_{\chi_{1}}}(\vec{\chi}_{1}\,)\otimes\psi_{n_{\chi_{2}}l_{\chi_{2}}}(\vec{\chi}_{2}\,)\otimes\psi_{n_{\chi_{3}}l_{\chi_{3}}}(\vec{\chi}_{3})$
where $\psi_{n_{\chi_{i}}l_{\chi_{i}}}$ are harmonic oscillator wave functions. The sum $n_{\chi_{i}},l_{\chi_{i}}$} is over $n_{\chi_{1}},n_{\chi_{2}},n_{\chi_{3}},l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}$. $N$ and $L$ are respectively the total principle quantum number and total angular momentum, and $L=l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$ while $l_{\chi_{1}},\,l_{\chi_{2}},\,l_{\chi_{3}}\leq 1$ for P wave tetraquark and $N=(2n_{\chi_{1}}+l_{\chi_{1}})+(2n_{\chi_{2}}+l_{\chi_{2}})+(2n_{\chi_{3}}+l_{\chi_{3}})$.
@@
Appendix B Cross term between different configuration from central potential
Table 9: Expectation values of the color-spin operator for the S=0 tetraquark state. $\psi^{k1}_{c}\psi^{h1}_{s}|\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}|\psi^{k2}_{c}\psi^{h2}_{s}>$
@@ Appendix B Cross term between different configuration from central potential
Table 10: Expectation values of the color-spin operator for S=2 tetraquark state. $\psi^{k1}_{c}\psi^{h1}_{s}|\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}|\psi^{k2}_{c}\psi^{h2}_{s}>$
@@
The color operator $\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}$ in the Hamiltonian in Eq. (II.2), along with the mass-dependent coefficient in the Cornell-like potential, mixes different color configurations, leading to nonzero cross terms for charmonium-like and bottomonium-like tetraquarks.
Similarly, the color-spin operator $\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}$, along with the mass-dependent factor in the spin-spin interaction,
@@
mixes different color-spin configurations, also leading to nonzero cross terms for charmonium-like and bottomonium-like tetraquarks.